# SeedVR2: One-Step Video Restoration via Diffusion Adversarial Post-Training

### v1

@@ 3.3 Training Procedures
Large-scale adversarial training is challenging. Benefiting from the low-quality condition in VR, we do not observe mode collapse [11] when starting from adversarial training. However, undesired artifacts can be observed after training for thousands of iterations, indicating that the unstable training issue still exists. Our approach improves the training stability from the following two aspects, i.e., distillation and loss.
Progressive Distillation. Directly adopting adversarial training to obtain a one-step model from an initial multi-step one may undermine the restoration ability of the model due to the large gap between the initial model and the target model. We conduct progressive distillation [53] to alleviate such a problem. To be specific, we start with the teacher model initialized from SeedVR [67] with $64$ sampling steps and progressively distill the student model to one step with a distillation stride of $2$. Each distillation procedure takes about $10$K iterations with a simple mean squared error loss. We also progressively increase the temporal length of the training data from images to video clips with a diverse number of frames during adversarial training, leading to robust VR performance toward videos with various lengths, including images. Benefiting from such a training strategy, we further obtain a 3B model distilled from the original 7B one, achieving comparable performance with only half of the model size.
Loss Improvement. Inspired by R3GAN [19], we first replace the non-saturating GAN loss [11] used in APT by a RpGAN loss [20] to avoid the potential mode dropping problem. We further introduce an approximate R2 regularization to penalize the gradient norm of $D$ on fake data while supporting modern deep learning software stacks:
@@
Besides GAN loss, L1 loss and LPIPS loss are commonly used in VR for perception-distortion tradeoff [2]. However, to compute LPIPS loss, we have to first decode the prediction from the latent space to pixel space, leading to an unaffordable computational cost in our scenario. Instead of LPIPS loss, we propose to adopt a feature matching loss via directly extracting features from the discriminator for efficient loss calculation. Specifically, we extract the features of predictions and ground-truths before the attention-only transformer blocks (i.e., the 16th, 26th, and 36th blocks of the transformer backbone) of the discriminator. Then, our feature matching loss $\mathcal{L}_{F}$ can be written as:
@@
where $D^{F}_{i}(\cdot)$ denotes the feature from the i-th block of discriminator. By default, we set the loss weight as $1.0$ for L1 loss, feature matching loss, and GAN loss when updating the generator. When updating the discriminator, we apply a weight of $1.0$ for GAN loss and the weights of the approximate R1 and R2 regularization are both $1000$. Note that the discriminator is fixed when updating the generator. In this way, the discriminator in our feature matching loss acts in a similar way to the VGG network [58] in LPIPS loss. Besides, the feature matching loss should also work with other GAN losses [11, 1, 45, 13] to further stabilize adversarial training for restoration tasks.
@@ 4 Experiments
Implementation Details. We train SeedVR2 on 72 NVIDIA H100-80G GPUs with around 100 frames of 720p per batch with sequence parallel [25] and data parallel [29]. Each stage of training takes about one day. We first train a 7B SeedVR model [67] from scratch following the new attention design in this paper. Then, we initialize the model parameters from 7B SeedVR model and follow the training strategies discussed in Sec. 3.3 for our SeedVR2 models. We mostly follow the training settings in APT [34] for adversarial training. We follow UAV [97] to synthesize about 10M image pairs and 5M video pairs for training.
Experimental Settings. Following previous work [97], we evaluate synthetic benchmarks, including SPMCS [80], UDM10 [64], REDS30 [46], and YouHQ40 [97], applying the same degradation settings as in training. The test resolution is 720p with an upscaling factor of $4$. Furthermore, we assess performance on the commonly used real-world dataset (VideoLQ [6]) and a self-collected AIGC dataset (AIGC28), which comprises 28 AI-generated videos with diverse resolutions and scenes. We employ a range of metrics to assess both frame-level and overall video quality. For synthetic pair datasets, we adopt full-reference metrics, including PSNR, SSIM, LPIPS [93], and DISTS [9]. For real-world and AI-generated content (AIGC) test data, where ground truth is unavailable, we rely exclusively on no-reference metrics, i.e., NIQE, CLIP-IQA, MUSIQ, and DOVER [^1] We adopt the technical score ranging from 0 to 100 following the official code.. To ensure test efficiency, the maximum output resolution is constrained to 1080p, with duration unchanged.
@@
Datasets Metrics RealViformer [95] MGLD-VSR [79] UAV [97] VEnhancer [14] STAR [74] SeedVR-7B [67] Ours 3B Ours 7B
SPMCS PSNR $\uparrow$ 24.185 23.41 21.69 18.20 22.58 20.78 22.97 22.90
@@ 4.1 Comparison with Existing Methods
Quantitative Comparisons. We compare our approach with all state-of-the-art real-world video restoration approaches. For diffusion-based methods [79, 97, 14, 74, 67], we adopt 50 sampling steps with a wavelet color fix post-processing [68], and keep other official settings unchanged. As shown in Table 1, our approach demonstrates superior performance in terms of perceptual metrics such as LPIPS and DISTS on synthetic benchmarks including SPMCS, UDM10 and YouHQ40. Note that RealViformer [95] and MGLD-VSR [79] involve REDS [46] in the train data, leading to high performance on the corresponding test set. As for real-world benchmarks, our method achieves comparable performance compared with other diffusion-based methods on VideoLQ and further obtains the highest NIQE, MUSIQ and DOVER scores on AIGC28, demonstrating our effectiveness.
Qualitative Comparisons. As observed in several previous studies [84, 86, 2, 12], existing image and video quality assessment metrics do not perfectly align with human perception. For example, non-reference metrics such as MUSIQ [23] and CLIP-IQA [66] prefer sharp results but may ignore the quality of details. We notice that such a phenomenon becomes more evident under high resolutions, e.g., 1080p. As shown in Figure 4.1, while our method does not show dominant metric performance on VideoLQ, the results generated by our approach are comparable to SeedVR [67] and outperform other baselines by a large margin.
@@
RealViformer-1 [95] +2% -38% -32%
VEnhancer-50 [14] -82% -86% -94%
UAV-50 [98] 0% -26% -26%
MGLD-VSR-50 [79] 0% -12% -12%
STAR-50 [74] +4% -22% -24%
SeedVR-7B-50 [67] +2% +10% +10%
@@
User Study. To further validation, we follow APT [34] to conduct a GSB test, i.e., the preference score is calculated as $\frac{G-B}{G+S+B}$, where G is the number of good samples preferred by the subjects, B is the bad samples not preferred, and S denotes the number of similar samples without preference. Thus, the score ranges from $-100\%$ to $100\%$ and $0\%$ indicates equal performance. We randomly select 25 samples from VideoLQ [6] and AIGC28, respectively, resulting in 50 LQ videos for test in total. We set our approach (7B) as the datum and compare it with existing methods [95, 14, 97, 79, 74, 67]. Given the LQ videos as reference, three experts are asked to evaluate the generated video quality from the following three criteria: visual fidelity, visual quality and overall quality. The visual fidelity measures the content similarity between the LQ reference and the generated result. The visual quality focuses on the realism of the generated results. The overall quality indicates the final preference after taking the above two factors into account. The subjects are given a pair of videos generated by different methods each time and asked to make their preferences for each criteria.
As shown in Table 2, our approach is comparable to the multi-step SeedVR [67] and clearly excels other approaches with better visual quality, aligning with the visual results shown in Figure 4.1. Particularly, VEnhancer [14] focuses on generative restoration, thus showing poor fidelity in real-world VR scenarios. Restricted by the limited generative capability of the diffusion prior, existing approaches [97, 79, 74] tend to generate inferior results with high-resolution inputs, indicating the necessity to train a large VR model without relying on the fixed prior. While our methods, i.e., ours-3B and ours-7B clearly outperform several baselines [95, 79, 97, 14, 74], the performance between these two models is different. Specifically, ours-3B receives more preference from the subjects than ours-7B, aligning with the results in Table 1. Recall that ours-3B is distilled from the 7B initial model. Such a performance gain may indicate the effectiveness of the distillation stage. And we believe our 7B model could receive further improvement with the scaling of computational resources.
[…]
figure Qualitative comparisons on both real-world [6] and AIGC videos. With a single sampling step, our SeedVR2 achieves comparable performance to SeedVR [67], and further excels other baselines with superior restoration capabilities, i.e., successfully removing the degradations while maintaining the textures of the bird, text, building, and the dog’s face (Zoom-in for best view).
@@ 4.2 Ablation Study
The Effect of Adaptive Window Attention. We first examine the effectiveness of the proposed adaptive window attention. We train the model with the predefined-size window attention and the proposed adaptive window attention, respectively. Both models share the same training settings for 20k iterations. As shown in Figure 4.2, when generating high-resolution (e.g., 1080p) results, window boundary inconsistency can be observed with a predefined-size attention window. We conjecture that such drawbacks indicate the limited model capability of handling overlapping windows, which is associated with the improper setting of window size compared with training resolutions. Specifically, applying a $64\times 64$ window over the compressed latent with a downsampling factor of $8$ makes the model insufficiently trained on window-overlapping cases, which are rare on the 720p training pairs. Moreover, we find that the diffusion transformer with RoPE embeddings [63] shows more robust performance across a range of resolutions after training on data with various sizes. Shifting to the window attention with mostly predefined window size [67] may weaken the generalization ability on other window sizes, i.e., the variable-sized windows near the boundary as shown in Figure 4.2. We show that the proposed adaptive window attention can significantly improve the model robustness by eliminating the above bad cases.
[…]
figure Comparisons of the window attention with a predefined size (i.e., ours w/ predefined win. atten.) and our adaptive window attention (i.e., ours w/ adaptive win. atten.). Boundary artifacts can be observed on high-resolution restoration with the predefined-size window attention (Zoom-in for best view).
Table 3: Ablation study on training losses and progressive training. For fairness, all baselines are trained on 72 NVIDIA H100-80G cards for 20k iterations. The comparison is conducted on YouHQ40 [97].
Metrics Non-satu. + R1 RpGAN + R1 + R2 RpGAN + R1 + R2 + L1 RpGAN + R1 + R2 + L1 + LF Prog. Training
@@
The Effect of Losses and Progressive Distillation. Training a large-scale GAN can be challenging due to its unstable nature. We verify the significance of various losses used in our method. We train each baseline with different loss combinations for 20k iterations and keep other settings the same. As shown in Table 3, compared with the vanilla loss used in APT [34] (i.e., non-saturating GAN loss [11] + R1), the model trained with RpGAN [20], R1 and R2 loss demonstrates significant improvement on perceptual metrics such as LPIPS and DISTS. We further observe a more stable training procedure without mode collapse, which exists under the settings of APT after long training. Besides, the adoption of L1 loss and the proposed feature matching loss both improve the metric performance, indicating the significance of these losses for restoration tasks. In practice, we notice that a large loss weight of L1 loss and feature matching loss improves the fidelity, but may lead to mildly over-smooth results compared with assigning a large weight to the GAN loss. Such an observation is consistent with the perception-distortion theory [2]. As a result, we reduce the loss weight of L1 loss and the feature matching loss to $0.1$ for the final model to enable better visual quality as reported in Sec. 4.1. Finally, as indicated in Table 3, applying a progressive distillation is necessary to maintain a strong restoration ability, which is expected since the distillation effectively minimizes the gap between the initial model and the one-step adversarial training.

### v2

@@ 3.3 Training Procedures
Large-scale adversarial training is challenging. Benefiting from the low-quality condition in VR, we do not observe mode collapse (Goodfellow et al., 2014) when starting from adversarial training. However, undesired artifacts can be observed after training for thousands of iterations, indicating that the unstable training issue still exists. Our approach improves the training stability from the following two aspects, i.e., distillation and loss.
Progressive Distillation. Directly adopting adversarial training to obtain a one-step model from an initial multi-step one may undermine the restoration ability of the model due to the large gap between the initial model and the target model. We conduct progressive distillation (Salimans and Ho, 2022) to alleviate such a problem. To be specific, we start with the teacher model initialized from SeedVR (Wang et al., 2025) with $64$ sampling steps and progressively distill the student model to one step with a distillation stride of $2$. Each distillation procedure takes about 1010K iterations with a simple mean squared error loss. We also progressively increase the temporal length of the training data from images to video clips with a diverse number of frames during adversarial training, leading to robust VR performance toward videos with various lengths, including images. Benefiting from such a training strategy, we further obtain a 3B model distilled from the original 7B one, achieving comparable performance with only half of the model size.
Loss Improvement. Inspired by R3GAN (Huang et al., 2024), we first replace the non-saturating GAN loss (Goodfellow et al., 2014) used in APT by a RpGAN loss (Jolicoeur-Martineau, 2019) to avoid the potential mode dropping problem. We further introduce an approximate R2 regularization to penalize the gradient norm of $D$ on fake data while supporting modern deep learning software stacks:
@@
Besides GAN loss, L1 loss and LPIPS loss are commonly used in VR for perception-distortion tradeoff (Blau and Michaeli, 2018). However, to compute LPIPS loss, we have to first decode the prediction from the latent space to pixel space, leading to an unaffordable computational cost in our scenario. Instead of LPIPS loss, we propose to adopt a feature matching loss via directly extracting features from the discriminator for efficient loss calculation. Specifically, we extract the features of predictions and ground-truths before the attention-only transformer blocks (i.e., the 16th, 26th, and 36th blocks of the transformer backbone) of the discriminator. Then, our feature matching loss $\mathcal{L}_{F}$ can be written as:
@@
where $D^{F}_{i}(\cdot)$ denotes the feature from the i-th block of the discriminator. By default, we set the loss weight as $1.0$ for L1 loss, feature matching loss, and GAN loss when updating the generator. When updating the discriminator, we apply a weight of $1.0$ for GAN loss and the weights of the approximate R1 and R2 regularization are both $1000$. Note that the discriminator is fixed when updating the generator. In this way, the discriminator in our feature matching loss acts in a similar way to the VGG network (Simonyan and Zisserman, 2015) in LPIPS loss. Besides, the feature matching loss should also work with other GAN losses (Goodfellow et al., 2014, Arjovsky et al., 2017, Mao et al., 2017, Gulrajani et al., 2017) to further stabilize adversarial training for restoration tasks.
@@ 4 Experiments
Implementation Details. We train SeedVR2 on 72 NVIDIA H100-80G GPUs with around 100 frames of 720p per batch with sequence parallel (Korthikanti et al., 2023) and data parallel (Li et al., 2020). Each stage of training takes about one day. We first train a 7B SeedVR model (Wang et al., 2025) from scratch following the new attention design in this paper. Then, we initialize the model parameters from 7B SeedVR model and follow the training strategies discussed in Sec. 3.3 for our SeedVR2 models. We mostly follow the training settings in APT (Lin et al., 2025) for adversarial training. We follow UAV (Zhou et al., 2024) to synthesize about 10M image pairs and 5M video pairs for training. During the distillation, loss is calculated on the vector field following Flow matching (Lipman et al., 2023). Both teacher and student models adopt the linear noise schedule with a timestep between 0 and 999. The teacher model uses the Euler sampler during training with a CFG scale of 7.5 for 64 timesteps and 1.0 for others. We adopt AdamW (Kingma and Ba, 2014) with a weight decay of 0.01 as optimizer, and the learning rate is set to $1\times 10^{-6}$.
Experimental Settings. Following previous work (Zhou et al., 2024), we evaluate synthetic benchmarks, including SPMCS (Yi et al., 2019), UDM10 (Tao et al., 2017), REDS30 (Nah et al., 2019), and YouHQ40 (Zhou et al., 2024), applying the same degradation settings as in training. The test resolution is 720p with an upscaling factor of $4$. Furthermore, we assess performance on the commonly used real-world dataset (VideoLQ (Chan et al., 2022a)) and a self-collected AIGC dataset (AIGC28), which comprises 28 AI-generated videos with diverse resolutions and scenes. We employ a range of metrics to assess both frame-level and overall video quality. For synthetic pair datasets, we adopt full-reference metrics, including PSNR, SSIM, LPIPS (Zhang et al., 2018b), and DISTS (Ding et al., 2020). For real-world and AI-generated content (AIGC) test data, where ground truth is unavailable, we rely exclusively on no-reference metrics, i.e., NIQE (Mittal et al., 2012), CLIP-IQA (Wang et al., 2023), MUSIQ (Ke et al., 2021), and DOVER (Wu et al., 2023) [^1] We adopt the technical score ranging from 0 to 100 following the official code.. To ensure test efficiency, the maximum output resolution is constrained to 1080p, with duration unchanged.
@@
Datasets Metrics RealViformer MGLD-VSR UAV VEnhancer STAR SeedVR-7B Ours-3B Ours-7B
SPMCS PSNR $\uparrow$ 24.19 23.41 21.69 18.20 22.58 20.78 22.97 22.90
@@ 4.1 Comparison with Existing Methods
Quantitative Comparisons. We compare our approach with all state-of-the-art real-world video restoration approaches. For diffusion-based methods, i.e., MGLD-VSR (Yang et al., 2024), UAV (Zhou et al., 2024), VEnhancer (He et al., 2024a), STAR (Xie et al., 2025), SeedVR-7B (Wang et al., 2025), we adopt 50 sampling steps with a wavelet color fix post-processing (Wang et al., 2024b), and keep other official settings unchanged. As shown in Table 1, our approach demonstrates superior performance in terms of perceptual metrics such as LPIPS and DISTS on synthetic benchmarks including SPMCS, UDM10 and YouHQ40. Note that RealViformer (Zhang and Yao, 2024) and MGLD-VSR involve REDS in the train data, leading to high performance on the corresponding test set. As for real-world benchmarks, our method achieves comparable performance compared with other diffusion-based methods on VideoLQ and further obtains the highest NIQE, MUSIQ and DOVER scores on AIGC28, demonstrating our effectiveness.
Qualitative Comparisons. As observed in several previous studies (Yu et al., 2024, Yue and Loy, 2024, Blau and Michaeli, 2018, Gu et al., 2022), existing image and video quality assessment metrics do not perfectly align with human perception. For example, non-reference metrics such as MUSIQ and CLIP-IQA prefer sharp results but may ignore the quality of details. We notice that such a phenomenon becomes more evident under high resolutions, e.g., 1080p. As shown in Figure 3, while our method does not show dominant metric performance on VideoLQ, the results generated by our approach are comparable to SeedVR and outperform other baselines by a large margin.
Figure 3: Qualitative comparisons on both real-world (Chan et al., 2022a) and AIGC videos. With a single sampling step, our method achieves comparable performance to SeedVR, further excelling other baselines with superior restoration capabilities, i.e., successfully removing the degradations while maintaining the textures of the bird, text, building, and the dog’s face (Zoom-in for best view).
User Study. To further validation, we follow APT (Lin et al., 2025) to conduct a GSB test, i.e., the preference score is calculated as $\frac{G-B}{G+S+B}$, where G is the number of good samples preferred by the subjects, B is the bad samples not preferred, and S denotes the number of similar samples without preference. Thus, the score ranges from $100\%$ to $100\%$ and $0\%$ indicates equal performance.
@@
RealViformer-1 +2% -38% -32%
VEnhancer-50 -82% -86% -94%
UAV-50 0% -26% -26%
MGLD-VSR-50 0% -12% -12%
STAR-50 +4% -22% -24%
SeedVR-7B-50 +2% +10% +10%
@@
We randomly select 25 samples from VideoLQ (Chan et al., 2022a) and AIGC28, respectively, resulting in 50 LQ videos for test in total. We set our approach (7B) as the datum and compare it with existing methods (Zhang and Yao, 2024, He et al., 2024a, Zhou et al., 2024, Yang et al., 2024, Xie et al., 2025, Wang et al., 2025). Given the LQ videos as reference, three experts are asked to evaluate the generated video quality from the following three criteria: visual fidelity, visual quality and overall quality. The visual fidelity measures the content similarity between the LQ reference and the output. The visual quality focuses on the realism of the generated results. The overall quality indicates the final preference after taking the above two factors as well as temporal consistency into account. The subjects are given a pair of videos generated by different methods each time and asked to make their preferences for each criterion.
As shown in Table 2, our approach is comparable to the multi-step SeedVR and clearly excels other approaches with better visual quality, aligning with the visual results shown in Figure 3. Particularly, VEnhancer focuses on generative restoration, thus showing poor fidelity in real-world VR scenarios. Restricted by the limited generative capability of the diffusion prior, existing approaches (Zhou et al., 2024, Yang et al., 2024, Xie et al., 2025) tend to generate inferior results with high-resolution inputs, indicating the necessity to train a large VR model without relying on the fixed prior. While our methods, i.e., ours-3B and ours-7B, clearly outperform several baselines (Zhang and Yao, 2024, Yang et al., 2024, Zhou et al., 2024, He et al., 2024a, Xie et al., 2025), the performance between these two models is different. Specifically, ours-3B receives more preference from the subjects than ours-7B, aligning with the results in Table 1. Recall that ours-3B is distilled from the 7B initial model. Such a performance gain may indicate the effectiveness of the distillation stage. And we believe our 7B model could receive further improvement with the scaling of computational resources.
@@ 4.2 Ablation Study
The Effect of Adaptive Window Attention. We first examine the effectiveness of the proposed adaptive window attention. We train the model with the predefined-size window attention and the proposed adaptive window attention, respectively. Both models share the same training settings for 20k iterations. As shown in Figure 4, when generating high-resolution (e.g., 1080p) results, window boundary inconsistency can be observed with a predefined-size attention window. We conjecture that such drawbacks indicate the limited model capability of handling overlapping windows, which is associated with the improper setting of the window size compared to the training resolutions. Specifically, applying a $64\times 64$ window over the compressed latent with a downsampling factor of $8$ makes the model insufficiently trained on window-overlapping cases, which are rare on the 720p training pairs. Moreover, we find that the diffusion transformer with RoPE embeddings (Su et al., 2024) shows more robust performance across a range of resolutions after training on data with various sizes. Shifting to the window attention with mostly predefined window size (Wang et al., 2025) may weaken the generalization ability on other window sizes, i.e., the variable-sized windows near the boundary as shown in Figure 4. We show that the proposed adaptive window attention significantly improves the model robustness by mitigating the aforementioned failure cases.
Figure 4: Comparisons of the window attention with a predefined size (i.e., ours w/ predefined win. atten.) and our adaptive window attention (i.e., ours w/ adaptive win. atten.). Boundary artifacts can be observed on high-resolution restoration with the predefined-size window attention.
Table 3: Ablation study on training losses and progressive training. All baselines are trained on 72 NVIDIA H100-80G cards for 20k iterations. The comparison is conducted on YouHQ40 (Zhou et al., 2024). Note that the first four baselines are trained w/o progressive training, while the last one is trained following our proposed method, but with different iterations for fair comparison.
Metrics Non-satu. + R1 RpGAN + R1 + R2 RpGAN + R1 + R2 + L1 RpGAN + R1 + R2 + L1 + LF w/ Prog. Training
@@
The Effect of Losses and Progressive Distillation. Training a large-scale GAN can be challenging due to its unstable nature. We verify the significance of various losses used in our method. We train each baseline with different loss combinations for 20k iterations and keep other settings the same. As shown in Table 3, compared with the vanilla loss used in APT (Lin et al., 2025) (i.e., non-saturating GAN loss (Goodfellow et al., 2014) + R1), the model trained with RpGAN (Jolicoeur-Martineau, 2019), R1 and R2 losses demonstrate significant improvement on perceptual metrics such as LPIPS and DISTS. We further observe a more stable training procedure without mode collapse, which exists under the settings of APT after long training. Besides, the adoption of L1 loss and the proposed feature matching loss both improve the metric performance, indicating the significance of these losses for restoration tasks. In practice, we notice that a large loss weight of L1 loss and feature matching loss improves the fidelity, but may lead to mildly over-smooth results compared with assigning a large weight to the GAN loss. Such an observation is consistent with the perception-distortion theory (Blau and Michaeli, 2018). As a result, we reduce the loss weight of L1 loss and the feature matching loss to $0.1$ for the final model to enable better visual quality as reported in Sec. 4.1. Finally, as indicated in Table 3, applying a progressive distillation is necessary to maintain a strong restoration ability, which is expected since the distillation effectively minimizes the gap between the initial model and the one-step adversarial training.
Figure 5: Qualitative results on real-world videos with challenging conditions, e.g., heavy degradations, large motion, and complex dynamics. Video comparison can be found in the demo video in the supplementary materials (Zoom-in for best view).
4.3 Limitations
The effectiveness of our SeedVR2 can be verified under challenging conditions as shown in Figure 5. However, we further identify several limitations of current SeedVR2 in practice. While our one-step method significantly saves time during sampling, the causal video VAE requires over 4x more time to encode and decode a video compared to the naive VAE commonly used by existing methods (Yang et al., 2024, Zhou et al., 2024, He et al., 2024a, Xie et al., 2025). In addition, when dealing with a 720p video with 100 frames, the casual video VAE takes over 95% of the total time. Enhancing the efficiency of the video VAE while maintaining performance is a worthwhile direction for future work.
Besides the VAE efficiency, we notice that our method is sometimes not robust to heavy degradations and very large motions, and shares some failure cases with existing methods, e.g., it may fail to fully remove degradations or generate visually unpleasing details. Moreover, due to the strong generation ability, SeedVR2 tends to overly generate details on inputs with very light degradations, e.g., 720p AIGC videos, leading to oversharpened results occasionally. Thus, we have to tune the model with careful hyperparameter settings. Improving the robustness towards complex real-world degradations and ensuring a satisfactory lower bound of performance remains a challenge for future work.